I have two points. (4,-6) and (-2,0)
how do I create two equations that represent the line that passes through these points?
Thank you

Answer by stanbon(75887)

I have two points. (4,-6) and (-2,0)
how do I create two equations that represent the line that passes through these points?
-----
Two points determine one line, not two.
-----
Form:: y = mx + b is the form of the equation of the line.
---
slope = m = (0--6)/(-2-4) = 6/-6 = -1
------
Solve for "b" using one of the points and the fact that m = -1
------
Using (-2,0) you get:
0 = -1*-2 + b
b = -2
------
Equation:
y = -x - 2
-----------
Cheers,
Stan H.
---------------